OpenStudy (yuii):

The equation of a line is 4x − 3y = −24 What is the x-intercept of the line? x =

OpenStudy (isaiahhlindsayy):

you are going to put into the (y=mx+b) formula

OpenStudy (welshfella):

the intercept occurs when y = 0 so plug in y = o and solve for x:- 4x - 3(0) = -24

OpenStudy (welshfella):

4x - 0 = -24 4x = -24 x = ?

OpenStudy (yuii):

-6.

OpenStudy (welshfella):

correct

OpenStudy (isaiahhlindsayy):

Yes

OpenStudy (yuii):

Thank you.